### v1

# Efficient Redundancy Reduction for Open-Vocabulary Semantic Segmentation

Abstract
Open-vocabulary semantic segmentation (OVSS) is an open-world task that aims to assign each pixel within an image to a specific class defined by arbitrary text descriptions. Recent advancements in large-scale vision-language models have demonstrated their open-vocabulary understanding capabilities, significantly facilitating the development of OVSS. However, most existing methods suffer from either suboptimal performance or long latency. This study introduces ERR-Seg, a novel framework that effectively reduces redundancy to balance accuracy and efficiency. ERR-Seg incorporates a training-free Channel Reduction Module (CRM) that leverages prior knowledge from vision-language models like CLIP to identify the most relevant classes while discarding others. Moreover, it incorporates Efficient Semantic Context Fusion (ESCF) with spatial-level and class-level sequence reduction strategies. CRM and ESCF result in substantial memory and computational savings without compromising accuracy. Additionally, recognizing the significance of hierarchical semantics extracted from middle-layer features for closed-set semantic segmentation, ERR-Seg introduces the Hierarchical Semantic Module (HSM) to exploit hierarchical semantics in the context of OVSS. Compared to previous state-of-the-art methods under the ADE20K-847 setting, ERR-Seg achieves +$5.6\%$ mIoU improvement and reduces latency by $67.3\%$. The project page is available at https://lchen1019.github.io/ERR-Seg.
Index Terms:
Open-vocabulary semantic segmentation, vision-language model, redundancy reduction
I Introduction
Semantic segmentation is a fundamental task in computer vision that aims to achieve pixel-level semantic understanding. Modern semantic segmentation methods [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] rely heavily on large-scale labeled datasets, but can only achieve understanding of limited predefined categories. In contrast, humans can understand scenes in an open-vocabulary manner, enabling the comprehension of arbitrary categories. The pursuit of open-vocabulary semantic segmentation (OVSS) is to replicate this intelligence that can perform unbounded pixel-level semantic understanding. Encouragingly, through natural language supervision, recent large-scale visual-language pre-training models (VLMs) exhibit open-vocabulary comprehension ability. However, existing VLMs receive image-level supervision, which is inadequate for the pixel-level representation required by semantic segmentation. This representation granularity gap impedes the transfer of open-world knowledge from VLMs to OVSS.
Figure 1: Performance vs. latency on ADE20K-847. Compared with ZegFormer [11], OVSeg [12], DeOP [13], SAN [14], SED [15] and CAT-Seg [16], ERR-Seg achieves a new state-of-the-art mIoU of $16.9\%$ with lower latency.
For this issue, recent efforts can be categorized into mask-based and cost-based methods. Mask-based methods [11, 17, 12, 18, 19, 14, 20, 21, 22, 23] model OVSS as the combination of mask generation and open-vocabulary mask classification, these approaches require region-level representation, which is much closer to the image-level representation. However, mask-based methods depend on a region proposal network (RPN), which is typically trained on a specific dataset, lacking the region proposal capability in the open-vocabulary context. To address this issue, some attempts [24, 25, 23, 26] employ the SAM [27] as a universal RPN. Nevertheless, as SAM tends to over-segment an instance into sub-patches, these methods have not made significant progress yet. In contrast, cost-based methods [16, 15] directly generate coarse-grained cost maps by calculating the cosine similarity between dense pixel-level embeddings and text embeddings. These methods effectively bridge the representation granularity gap and eliminate the requirement of RPN. As a result, the methods in this family have shown superior performance compared to mask-based methods.
Although cost-based methods have effectively addressed the challenge of representation granularity, there are still some issues to be further considered: (1) Redundancy: It is evident that real-world categories are limitless, while those depicted in an image are always limited, leading to numerous redundant classes. OVSS tasks struggle with this redundancy. On the one hand, it leads to increased latency due to the redundant channel dimension of the cost maps. On the other hand, it poses a challenge for the attention mechanism to model class-level sequences containing numerous redundant classes. (2) Hierarchical Semantic: Existing closed-set semantic segmentation methods [28, 2] typically extract hierarchical semantic maps from middle-layer features and aggregate them to derive the final semantic maps. In contrast, cost-based methods only integrate semantic information from pixel-text cost maps, neglecting the semantic details in middle-layer features.
To this end, we propose ERR-Seg, a novel cost-based framework for OVSS that jointly considers efficiency and accuracy. To address the issue (1), ERR-Seg presents the Channel Reduction Module (CRM) to select the most relevant categories while discarding others based on the category existence probability implied in the cost maps. By reducing the channel dimensions, the attention mechanism can better capture spatial-level long-range dependencies, as depicted in Fig. 2. This reduction also significantly decreases latency. Furthermore, ERR-Seg introduces Efficient Semantic Context Fusion (ESCF) with Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR), which reduces computational workload by $49.3\%$ without compromising performance. To address the issue (2), ERR-Seg incorporates the Hierarchical Semantic Module (HSM) to transform middle-layer features into semantic space to compute cost maps in parallel. These hierarchical cost maps are then merged by using a pyramid structure to aggregate more middle-layer semantic details. As shown in Fig. 1, ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy.
The main contributions can be summarized as follows:
  • We propose a training-free Channel Reduction Module (CRM) to eliminate redundant classes. By customizing a small-scale vocabulary list for each image, our proposed ERR-Seg can handle an extensive number of classes while significantly reducing computational complexity.
  • We quantitatively and qualitatively explore the impact of redundant classes on spatial-level and class-level contextual information modeling in the context of OVSS.
  • We present the Hierarchical Semantic Module (HSM) to generate hierarchical cost maps, enabling the aggregation of semantic details from middle-layer features.
  • We introduce an efficient framework for OVSS, which achieves superior performance on multiple datasets and significantly accelerates the inference process.
II Related Work
II-A Closed-set Semantic Segmentation
From the perspective of modeling semantic segmentation tasks, technical methods can be divided into two categories: (1) FCN-based methods [29] consider the semantic segmentation task as a dense pixel-by-pixel prediction task. Various FCN-based variants have been developed to incorporate more contextual information, such as dilated convolution [1, 7], feature pyramid [30, 28], pyramid pooling module [31, 28], object contextual representation [8], attention mechanisms [5, 32]. Additionally, some methods [6, 33, 2, 34, 9, 10] introduce transformer-based architectures to semantic segmentation. (2) Mask-based methods [35, 36, 3, 4] treat the segmentation task as a combination of mask generation and mask classification. These methods aim to unify various segmentation tasks (such as instance segmentation, semantic segmentation, and panoptic segmentation) and have demonstrated their effectiveness in the context of closed-set semantic segmentation.
Figure 2: Visual comparison of attention maps. With our proposed channel reduction module (CRM), the attention mechanism can more effectively capture spatial-level long-range dependencies in the context of OVSS.
Figure 3: Overall architecture of ERR-Seg. We first utilize the Hierarchical Semantic Module to generate dense pixel-text cost maps. The training-free Channel Reduction Module is then used to eliminate redundant classes. The Efficient Semantic Context Fusion further enhances the cost maps, improving spatial and class-level contextual information. Finally, the Gradual Upsampling Decoder restores the high-rank information of cost maps by incorporating image details from the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ of CLIP’s visual encoder.
II-B Vision-Language Models
Visual-language models (VLMs) aim to extend open-world comprehension to the visual domain by jointly training visual and language modalities. Early VLMs [37, 38, 39, 40] were trained on small-scale datasets through multiple multi-modal tasks. In contrast, CLIP [41] collects large-scale image-text pairs and uses image-text contrastive learning to align the visual and language modality. To further enhance the quality and efficiency of image-text pre-training, ALIGN [42] collected a much larger dataset of image-text pairs without removing noisy data, SigLIP [43] introduces pairwise sigmoid loss to eliminate the requirement of a global view of the pairwise similarities for normalization. Beyond image-text contrast learning, BLIP-2 [44] and LLaVA [45] introduce an additional language model to process aligned images and text in a unified framework for more complex reasoning. In this study, we further explore the use of prior knowledge from VLMs to better support the OVSS task.
II-C Open-Vocabulary Semantic Segmentation
Earlier methods [46, 47, 48] for OVSS attempt to establish a joint embedding space connecting image pixels and a predefined vocabulary list. However, these methods struggle with arbitrary class recognition due to their limited vocabulary size. Fortunately, vision-language models (VLM) [41, 42, 49] pre-trained on extensive web data exhibit strong capabilities in open-vocabulary recognition, leading recent research focused on adapting these models for OVSS. Based on the method of introducing semantic information from VLM, current OVSS methods can be categorized into two main groups: mask-based methods and cost-based methods.
For mask-based methods, an intuitive two-stage approach is first proposed. Initially, an additional RPN is adopted to propose region candidates, followed by the application of VLMs for open-vocabulary recognition. ZegFormer [11] and SimBaseline [17] directly adopt this two-stage method, leading to suboptimal performance and significant computational overload. To improve the performance, ODISE [19] utilizes a text-to-image diffusion model as RPN to generate high-quality masks, and OVSeg [12] fine-tunes CLIP on region-text pairs for improved open-vocabulary recognition. To further reduce the computational overload, single-stage methods have been introduced to avoid multiple forward propagations. Among these methods, MaskCLIP [50] and SAN [14] adopt attention bias, while FC-CLIP [20] employs mask pooling. To further enhance the performance of single-state methods, EBSeg [21] alleviates overfitting to training classes through image embedding balancing, and MROVSeg [22] introduces a multi-resolution training framework to tackle the problem that VLMs are typically pre-trained with low-resolution images.
For cost-based methods, CAT-Seg [16] aggregates the pixel-text cost maps at both spatial-level and class-level at one resolution. It also emphasizes the importance of fine-tuning the text encoder of CLIP. In contrast, SED [15] gradually aggregates the cost maps at different resolutions. However, cost-based methods have a drawback compared to mask-based methods: they involve operators with quadratic complexity in relation to the vocabulary size. This means that when the vocabulary is large, it can lead to significant computational and storage overhead. Our proposed ERR-Seg is also cost-based, but we introduce a training-free mechanism to remove redundant classes, thereby tailoring a small-scale vocabulary list for each image.
III Methodology
III-A Overall Architecture
The overview of our proposed ERR-Seg is shown in Fig. 3. Given an image $I\in\mathbb{R}^{3\times H\times W}$ and an arbitrary set of classes $\mathcal{C}=\{C_{1},...,C_{K}\}$, the OVSS task aims to assign a specific class $C_{i}$ to each pixel in $I$. Here, $K$ is the number of classes. Our framework comprises four stages: (1) Encoding the image $I$ to derive the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, as well as dense image embeddings $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$. Encoding the classes $\mathcal{C}$ to generate text embeddings of each class $\mathcal{T}$. (2) Generating the pixel-text cost maps by using $\mathcal{F}_{3}^{s}$, $\mathcal{F}_{4}^{s}$ and $\mathcal{T}$, and removing redundant classes to customize small-scale cost maps for each image. (3) Enhancing the cost maps by incorporating more spatial-level and class-level contextual information. (4) Leveraging the high-rank textual information from $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, to generate a more detailed high-resolution segmentation result.
III-B Hierarchical Semantic Module
Typically, cost-based methods introduce semantic information by calculating pixel-text cost maps using only the last layer output, which ignores semantic details from middle-layer features. To address this, we propose the Hierarchical Semantic Module (HSM) to extract hierarchical semantic details from the middle-layer features of open-vocabulary visual encoders. HSM calculates pixel-text cost maps on the middle-layer features in parallel and merges these hierarchical cost maps through a feature pyramid strategy.
Initially, the HSM first encodes $I$ to derive the middle-layer features $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$, and dense image-text aligned embeddings $\mathcal{F}_{4}^{s}$. The relationship between $\mathcal{F}_{3}^{v}$ and $\mathcal{F}_{4}^{s}$ is modeled as follows:
$\mathcal{F}_{4}^{s}$ $=g\left(f\left(conv_{s=2}(\mathcal{F}_{3}^{v})\right)\right),$ (1)
where $f$ represents stage 4 of CLIP’s visual encoder, $g$ is the linear projection head, and $conv_{s=2}$ denotes convolution with a stride of 2 for downsampling. Additionally, the HSM generates extra image-text aligned embeddings from $\mathcal{F}_{3}^{v}$:
$\mathcal{F}_{3}^{s}=g\left(f\left(conv_{s=1}(\mathcal{F}_{3}^{v})\right)\right),$ (2)
where $conv_{s=1}$ denotes convolution with a stride of 1, using the same initial values as $conv_{s=2}$. $conv_{s=1}$ prevents downsampling, thereby keeping the size of $\mathcal{F}_{3}^{s}$ the same as $\mathcal{F}_{3}^{v}$.
Figure 4: Pipeline of our proposed CRM mechanism. CRM involves a training-free scoring function to assign scores to each class, retaining the top-$P$ classes while eliminating other redundant classes.
Subsequently, HSM computes the hierarchical cost maps. Given a vocabulary list $\mathcal{C}=\{C_{1},...,C_{K}\}$, a set of sentences $\mathcal{G}=\{G_{1},...,G_{K}\}$ is first generated, where $G_{i}$ consists of $C_{i}$ and a well-crafted prompt (e.g., “a photo of $\{C_{i}\}$ in the scene”). $\mathcal{G}$ is then fed into the CLIP’s text encoder to get text embeddings for each class, denoted as $\mathcal{T}=\{T_{1},...,T_{K}\}$. The pixel-text cost maps are derived by calculating the cosine similarity between each pixel and each class:
$S(i,j,t)=\frac{\mathcal{F}^{s}(i,j)\cdot\mathcal{T}(t)}{\left\|\mathcal{F}^{s}(i,j)\right\|\left\|\mathcal{T}\left(t\right)\right\|},$ (3)
where $(i,j)$ denotes the 2D spatial positions, and $t$ represents the index of text embeddings. The $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ are used to derive pixel-text cost maps $S_{3}$ and $S_{4}$ individually.
To aggregate the hierarchical cost maps, we employ a feature pyramid strategy. More specifically,
$S=conv\left(S_{3}\right)+up\left(conv\left(S_{4}\right)\right),$ (4)
where $S_{3}$ and $S_{4}$ are first projected into the same dimension by $conv$, then $S_{4}$ is upsampled to align with the size of $S_{3}$ via $up$, and subsequently add them to yield final cost maps $S\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$.
III-C Channel Reduction Module
The Channel Reduction Module (CRM) is proposed to eliminate redundant classes, based on the observation that the number of classes within an image is limited. As illustrated in Table I, images annotated with $847$ classes have an average of only $10.2$ categories, indicating that there are $98.8\%$ redundant classes. Even images annotated with 20 classes have more than $90.0\%$ redundant classes. The impact of redundant classes will be discussed later in this section. CRM introduces a training-free channel selection function $\Phi$, which leverages the prior knowledge in original cost maps $S\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$ to convert $S$ into reduced cost maps $S_{r}=\Phi\left(S\right)$, where $S_{r}\in\mathbb{R}^{K\times C\times H_{c}\times W_{c}}$ and $K>P$. $P$ is an empirical value and will be discussed later.
TABLE I: Analysis of image category count.
PAS-20 PC-59 A-150 PC-459 A-847
Average 1.5 4.8 8.5 6.5 10.2
Redundancy $92.5\%$ $91.9\%$ $94.3\%$ $98.6\%$ $98.8\%$
The pipeline of the function $\Phi$ is illustrated in Fig. 4. Initially, $S$ is averaged across the $C$ dimension, resulting in $D\in\mathbb{R}^{K\times H_{c}\times W_{c}}$, representing the probability of each category at each spatial location. Subsequently, for each spatial location, the top $k$ classes are selected based on their probability values and arranged in descending order:
$N=Top(D,k),$ (5)
where $N\in\mathbb{R}^{k\times H_{c}\times W_{c}}$ represents the indices of the top $k$ categories. Then we calculate the scores for each category. The score for $t$-th category, denoted as $R_{t}$, is computed as:
$R_{t}=\sum_{c=1}^{k}{\lambda^{c-1}\left(\sum_{i=1}^{H_{c}}{\sum_{j=1}^{W_{c}}{\left[N_{c}(i,j)=t\right]}}\right)},$ (6)
where $\lambda^{c-1}$ represents the score assigned to the class ranked $c$. $\lambda$ is the score coefficient and is expected to meet the condition $10<\lambda<$ 1, as higher-ranked categories should receive higher scores. The notation []delimited-[][$\cdot$][ ⋅ ] represents a logical expression that yields 1 when the condition is true and 0 otherwise. After scoring all categories, we obtain =$\mathcal{R}=\left\{R_{i}\right\}_{i=1}^{K}$. Finally, $\mathcal{R}$ is sorted in descending order, and only the top $P$ categories are retained. Subsequently, $S_{r}$ is utilized instead of $S$ in the following calculations.
The original cost maps $S$ comprise the existing and redundant classes. Let $\mathcal{Q}_{e}=\{S(i)\}_{i=1}^{M}$ collect the original cost maps corresponding to the existing classes, where $M$ is the number of existing classes in $I$. The cost maps corresponding to the redundant classes are represented by $\mathcal{Q}_{r}=\{S(i)\}_{i=M+1}^{K}$, and those corresponding to the redundant classes after partial reduction are denoted as $\mathcal{Q}_{p}=\{S(i)\}_{i=M+1}^{P}$. In the subsequent subsections, we will analyze the impact of $\mathcal{Q}_{r}$ and $\mathcal{Q}_{p}$ on class-level and spatial-level attention mechanisms (the primary operations after CRM), as well as on latency.
III-C1 Impact on class-level attention.
The class-level attention is the attention mechanism applied to the class dimension of $S$, which is used in CFM as visualized in Fig. 3(b). Its output of the $q$-th class in $\mathcal{Q}_{e}$ before reduction is:
$O^{e,r}(q)=\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)},$ (7)
where $q\leq M$, and $A\in\mathbb{R}^{K\times K}$ stands for the class-level attention score matrix before reduction. The contribution of $\mathcal{Q}_{r}$ to the attention output of the $q$-th class is formulated as:
$O^{r}(q)=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)}.$ (8)
Similarly, we denote the output of the $q$-th class after reduction as $O^{e,p}(q)$, and the contribution of $\mathcal{Q}_{p}$ as $O^{p}(q)$.
Proposition 1.
The relationship between the contribution ratio of $\mathcal{Q}_{r}$, denoted as $\Delta_{r}$, and the contribution ratio of $\mathcal{Q}_{p}$, denoted as $\Delta_{p}$, is expressed as:
$\Delta_{p}=\frac{O^{p}(q)}{O^{e,p}(q)}<\Delta_{r}=\frac{O^{r}(q)}{O^{e,r}(q)}$ (9)
Proof.
By utilizing Eq. 7 and Eq. 8, the expression $\Delta r=O^{r}(q)/O^{e,r}(q)$ can be simplified as:
$\Delta_{r}=\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$ (10)
Similarly, $\Delta_{p}$ can be simplified as:
$\Delta_{p}=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)},$ (11)
where $B\in\mathbb{R}^{P\times P}$ represents the attention scores matrix after redundancy reduction. Let us define:
$\Delta_{r}^{\prime}=\left(\frac{1}{\Delta_{r}}-1\right),$ (12)
$\Delta_{p}^{\prime}=\left(\frac{1}{\Delta_{p}}-1\right).$ (13)
Since $10<\Delta_{r},\Delta_{p}<$ 1, it follows that $\Delta_{r}^{\prime},\Delta_{p}^{\prime}>0$. To compare $\Delta_{r}$ and $\Delta_{p}$, we first compare $\Delta_{r}^{\prime}$ and $\Delta_{p}^{\prime}$, by analyzing the following formula:
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\frac{\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}}\frac{\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}{\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}}.$ (14)
Let us define:
$\alpha=\sum_{k=1}^{M}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=1}^{M}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}$ (15)
and
$\beta=\sum_{k=M+1}^{P}{\exp\left(\frac{B_{qk}}{\sqrt{C}}\right)S(k)}/\sum_{k=M+1}^{K}{\exp\left(\frac{A_{qk}}{\sqrt{C}}\right)S(k)}.$ (16)
Thus, we have
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta.$ (17)
Since $A$ and $B$ arise from similarity computations between the same query and key, and redundant categories do not impact the similarity scores of common categories, we have
$\forall 1\leq i,j\leq P,A_{ij}=B_{ij}.$ (18)
This implies that $\alpha=1$ and $\beta<1$. Hence,
$\frac{\Delta_{r}^{\prime}}{\Delta_{p}^{\prime}}=\alpha\beta<1.$ (19)
It indicates that $\Delta_{r}^{\prime}<\Delta_{p}^{\prime}$. Thus, we can get $\Delta_{p}<\Delta_{r}$. In this way, we finish the proof. ∎
The above inequality suggests that as the number of redundant classes increases, they contribute more to the final output. Thus, existing classes contribute less. As such, we can deduce that reducing redundancy facilitates the class-level attention to focus more on existing classes.
Figure 5: Architecture of our proposed SFM and CFM. SSR signifies Spatial-level Sequence Reduction, while CSR represents Class-level Sequence Reduction. DW-Conv refers to deep-wise convolution, and FC indicates fully connected layers.
III-C2 Impact on spatial-level attention
Spatial-level attention aims to capture dependencies among different pixels within the same class, which is used in SFM as visualized in Fig. 3(a). Typically, it assigns high scores when two pixels are close or distant but belong to the same class, indicating similar embeddings. However, after class-level attention, each pixel-level embedding integrates contributions from redundant classes. This leads to decreased similarity among pixel-level embeddings of the same class, hindering spatial-level attention from effectively capturing dependencies on distant pixels within the same class. In Fig. 2, we visualize the spatial-level attention maps of different classes, confirming our conclusion that redundancy causes spatial-level attention to focus more on local regions and compromises the ability to capture long-range dependencies.
III-C3 Impact on latency
Without using CRM, the computational complexity of class-level and spatial-level attention is $O(K^{2}CH_{c}W_{c})$ and $O(KCH_{c}^{2}W_{c}^{2})$, respectively. With the help of CRM, the corresponding complexities become $O(P^{2}CH_{c}W_{c})$ and $O(PCH_{c}^{2}W_{c}^{2})$, which means that CRM can reduce the computation of approximately $1-\left({P}/{K}\right)^{2}$ and $1-{P}/{K}$ for class-level and spatial-level attention, respectively.
TABLE II: Comparison with state-of-the-art methods. Img. and Txt. indicate whether the VLM’s image and text encoder are fine-tuned, respectively. A.B. signifies whether an additional backbone is employed.
Method VLM Img. Txt. A.B. Training dataset A-847 PC-459 A-150 PC-59 PAS-20
ZegFormer [11] [CVPR’22] ViT-B/16 COCO-Stuff 5.6 10.4 18.0 45.5 89.5
SimBaseline [17] [ECCV’22] ViT-B/16 COCO-Stuff 7.0 - 20.5 47.7 88.4
OVSeg [12] [CVPR’23] ViT-B/16 COCO-Stuff 7.1 11.0 24.8 53.3 92.6
DeOP [13] [ICCV’23] ViT-B/16 COCO-Stuff-156 7.1 9.4 22.9 48.8 91.7
SAN [14] [CVPR’23] ViT-B/16 COCO-Stuff 10.1 12.6 27.5 53.8 94.0
EBSeg [21] [CVPR’24] ViT-B/16 COCO-Stuff 11.1 17.3 30.0 56.7 94.6
SED [15] [CVPR’24] ConvNeXt-B COCO-Stuff 11.4 18.6 31.6 57.3 94.4
CAT-Seg [16] [CVPR’24] ViT-B/16 COCO-Stuff 12.0 19.0 31.8 57.5 94.6
ERR-Seg [ours] ConvNeXt-B COCO-Stuff 11.7 18.9 32.9 56.6 94.9
ERR-Seg [ours] ConvNeXt-B COCO-Stuff 13.9 20.5 35.8 56.9 95.4
OVSeg [12] [CVPR’23] ViT-L/14 COCO-Stuff 9.0 12.4 29.6 55.7 94.5
ODISE [19] [CVPR’23] ViT-L/14 COCO Panoptic 11.1 14.5 29.9 57.3 -
SAN [14] [CVPR’23] ViT-L/14 COCO-Stuff 12.4 15.7 32.1 57.7 94.6
FC-CLIP [20] [NeurIPS’23] ConvNeXt-L COCO Panoptic 14.8 18.2 34.1 58.4 95.4
SAM-CP [23] [ArXiv’24] ConvNext-L COCO-Panoptic - - 31.8 - -
Unpair-Seg [51] [ArXiv’24] ConvNext-L Merged-130K 14.6 19.5 32.6 52.2 93.0
EBSeg [21] [CVPR’24] ViT-L/14 COCO-Stuff 13.7 21.0 32.8 60.2 96.4
SED [15] [CVPR’24] ConvNeXt-L COCO-Stuff 13.9 22.6 35.2 60.6 96.1
CAT-Seg [16] [CVPR’24] ViT-L/14 COCO-Stuff 16.0 23.8 37.9 63.3 97.0
ERR-Seg [ours] ConvNeXt-L COCO-Stuff 15.8 22.7 36.5 60.7 96.6
ERR-Seg [ours] ConvNeXt-L COCO-Stuff 16.9 23.9 38.3 60.7 96.6
TABLE III: Efficiency comparison. The latency (ms) is computed on a single NVIDIA RTX 3090 GPU.
Method A-847 A-150
Latency GFLOPs mIoU Latency GFLOPs mIoU
ZegFormer 158 1954 4.9 154 1932 16.9
SimBaseline 331 1878 7.0 329 1783 20.5
DeOP 503 2791 7.1 178 711 22.9
OVSeg 331 1791 7.1 324 1783 24.8
SAN 34 120 10.1 32 95 27.5
SED 196 571 11.4 114 291 31.6
CAT-Seg 552 2154 12.0 369 1516 31.8
ERR-Seg 114 348 13.9 85 288 35.8
III-D Efficient Semantic Context Fusion
It is well known that context is essential for semantic understanding [5, 32, 8]. Hence, ERR-Seg incorporates the Efficient Semantic Context Fusion (ESCF) to integrate spatial-level and class-level contextual information via the Spatial-level Fusion Module (SFM) and Class-level Fusion Module (CFM), respectively. Additionally, Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR) strategies are introduced to accelerate the context fusion process without compromising performance.
III-D1 Spatial-level Fusion Module
Spatial-level Fusion Model (SFM) aims to model the relationship between different spatial locations within the same class. The architecture of SFM is presented in Fig. 5(a). Inspired by [52], SFM introduces Spatial-level Sequence Reduction (SSR) to reduce the computational complexity of the attention mechanism. Specifically, it reduces the sequence length of $K$ and $V$ by a factor of $R$, as follows:
$Z^{\prime}$ $=\mathrm{Reshape}\left(Z,\left(N/R,C\times R\right)\right),$ (20)
$Z$ $=\mathrm{Linear}\left(Z^{\prime},\left(C\times R,R\right)\right),$ (21)
where $Z\in\{K,V\}$, $N$ is the sequence length, $\mathrm{Reshape(.)}$ refers to reshaping $Z$ into the shape of $\left(N/R,C\times R\right)$, and $\mathrm{Linear(.)}$ represents a fully connected layer that projects the input from $C\times R$ dimension to $C$ dimension. As a result, it merges adjacent $R$ embeddings into one embedding and reduces the complexity of the self-attention mechanism from $O(N^{2})$ to $O(N^{2}/R)$. The cost maps $S_{r}$ will then go through a depth-wise convolution layer, followed by a two-layer MLP. Moreover, we replace the first linear layer of the MLP with a star operation [53] to leverage its ability to map to high-dimensional spaces. More specifically, the star operation is defined as follows:
$S_{r}^{(1)}$ $=\mathrm{Linear}\left(S_{r},\left(C,C^{\prime}\right)\right),$ (22)
$S_{r}^{(2)}$ $=\mathrm{Linear}\left(S_{r},\left(C,C^{\prime}\right)\right),$ (23)
$S_{r}$ $=\sigma\left(S_{r}^{(1)}\right)*S_{r}^{(2)},$ (24)
where $\sigma$ denotes the activation function, and $C^{\prime}$ is the intermediate embedding dimension of the MLP.
III-D2 Class-level Fusion Module
Class-level fusion module (CFM) aims to capture dependencies within the same spatial location but across different classes. The architecture of our proposed CFM is depicted in Fig. 5(b). Noticing that surrounding pixels tend to exhibit similar features, CFM introduces the Class-level Sequence Reduction (CSR), which employs the average pooling to reduce the size of the cost maps and residual connections to incorporate the upsampling results and origin input.
III-E Gradual Upsampling Decoder
After the ESCF, low-resolution semantic maps $E_{r}\in\mathbb{R}^{P\times C\times H_{c}\times W_{c}}$ are obtained. To recover high-resolution semantic maps, we propose the Gradual Upsampling Decoder (GUD), leveraging the high-rank details from $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ to guide the upsampling process of $E_{r}$. The architecture of GUD is illustrated in Fig. 3. Initially, $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ undergoes a $3\times 3$ convolution and is then concatenated with the cost maps at each class. The concatenated features are then passed through the guidance refine block, comprising two $3\times 3$ convolutions. Our observations indicate that a small kernel yields better performance, as the texture is local and does not require a large receptive field. Moreover, we emphasize the significance of stopping the gradient of $\mathcal{F}_{i}^{v},i\in\{1,2,3\}$ as it can prevent the visual encoder from distribution shift.
IV Experiments
TABLE IV: Label-level similarity between evaluation and training datasets.
Dataset $1.00$ $\geq 0.95$ $\geq 0.90$
PAS-20 0.45 0.90 1.00
PC-59 0.30 0.85 0.95
A-150 0.19 0.41 0.57
PC-459 0.10 0.21 0.29
A-847 0.06 0.12 0.21
TABLE V: Component Analysis. We remove modules from ERR-Seg independently to evaluate their impact.
Module A-847 PC-459 A-150 PC-59 PAS-20
ERR-Seg 13.9 20.5 35.8 56.9 95.4
w/o HSM 13.3 19.3 35.5 56.4 95.0
w/o CRM 14.0 20.1 35.3 56.8 94.9
w/o ESCF 9.3 12.8 26.1 43.6 90.8
w/o GUD 13.4 19.7 34.8 56.4 95.1
TABLE VI: Ablation study for HSM. Different gradient sources of updating stage 4 of CLIP’s visual encoder are tried.
Stage 4 gradient from A-847 PC-459 A-150 PC-59 PAS-20
$\mathcal{F}_{3}^{s}$ 7.3 13.5 20.1 45.5 90.6
$\mathcal{F}_{4}^{s}$ 14.1 20.4 35.6 57.0 94.7
$\mathcal{F}_{3}^{s}+\mathcal{F}_{4}^{s}$ 13.9 20.5 35.8 56.9 95.4
IV-A Experimental Setup
IV-A1 Traning dataset
Following the existing OVSS methods [16, 15], we train our model on COCO-Stuff [54]. It has 118,287 images for training with 171 different classes.
IV-A2 Evaluation dataset
  • PASCAL-Context [55] contains 459 classes with 5,104 images for validation. We report the scores of the full classes (PC-459) and the scores of the 59 most frequent classes (PC-59).
  • ADE20K [56] has 2,000 validation images with 847 classes. Similarly to PASCAL-Context, we report the results of 150 frequent classes (A-150) and full 847 classes (A-847).
  • PASCAL VOC [57] includes 1,449 images for validation with 20 common object classes. We report the performance of the 20 classes named PAS-20.
IV-A3 Analysis of the evaluation dataset
Prior studies have typically overlooked the label-level similarity between the validation and training sets, while it is necessary for open vocabulary-related tasks. In this section, we extract text embeddings for each category and then calculate the cosine similarity between the categories in the evaluation and training datasets, resulting in a similarity matrix. Subsequently, we map each category in the evaluation dataset to a category with the highest similarity score in the training dataset. The results are presented in Table IV; we report the proportion of labels in the validation sets with a similarity score equal to $1.00$, greater than $0.95$, $0.90$, and $0.85$.
Most categories in PAS-20 and PC-59 display similarity scores exceeding $0.90$, indicating that a significant number of categories in these datasets were encountered during the training process. This also implies that overfitting to COCO-Stuff can lead to performance improvement in PAS-20 and PC-59. Hence, PAS-20 and PC-59 measure more in-domain open-vocabulary semantic segmentation capability. In contrast, the ratio is comparatively low for A-150, PC-459, and A-847, suggesting that these three datasets are suitable for measuring out-domain open-vocabulary capability.
IV-A4 Implementation details
We adopt the Convolutional CLIP [58] from OpenCLIP [59], where the visual encoder is ConvNeXt [60] architecture. Our code is implemented on the top of Detectron2 [61]. The AdamW [62] optimizer is employed, with an initial learning rate of $2\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. The learning rate is set to $2\times 10^{-6}$ for fine-tuning the CLIP model. Our model is trained on COCO-Stuff using cross-entropy loss for $80$K iterations, with a batch size of $4$ and a crop size of $640\times 640$. $4$ NVIDIA RTX 3090 GPUs were used for training.
Figure 6: Impact of score coefficient $\lambda$ on CRM.
Figure 7: Analysis of selected class number $P$ during training. It indicates that excessively large or small values of $P$ result in performance degradation.
IV-B Comparing with State-of-the-art
We compare our proposed ERR-Seg with existing state-of-the-art methods on the five standard evaluation settings, with the results presented in Table II. To ensure a fair comparison, all models utilize CLIP as the VLM. We also include whether CLIP’s image and text encoders are fine-tuned. When both the image and text encoders are fine-tuned, ERR-Seg achieves remarkable performance on these five settings, particularly in the base model. Notably, on A-847, PC-459, A-150 and PAS-20, ERR-Seg exhibited significant improvements, outpacing CAT-Seg [16] by $15.8\%$, $7.9\%$, $12.6\%$ and $0.8\%$, respectively. In the large model, ERR-Seg surpasses the previous arts on A-847, PC-459 and A-150 by $5.6\%$, $0.4\%$ and $1.1\%$, respectively. When only the text encoder is fine-tuned, ERR-Seg also outperforms SED [15] on all five settings, with an improvement of $13.7\%$, $0.4\%$, $5.2\%$, $0.2\%$ and $0.6\%$ on A-847, PC-459, A-150, PC-59 and PAS-20, respectively.
However, ERR-Seg does not perform as well as CAT-Seg on PC-59, exhibiting a $4.1\%$ lower performance than CAT-Seg in the large model. As discussed in Section IV-A3, PC-59 primarily assesses in-domain open-vocabulary performance. This suggests that CAT-Seg might be overfitting to the seen categories during training, thus showcasing superior in-domain performance compared to our model but lagging in out-domain performance. In contrast, with the introduction of CRM, ERR-Seg customizes the vocabulary list for each image during training, which helps prevent overfitting to the seen categories.
Table III presents the results of the latency comparisons. Although cost-based CAT-Seg shows notable performance enhancement over mask-based SAN, it also results in a considerable increase in latency. In contrast, our proposed ERR-Seg not only further improves the performance of cost-based methods, but also reduces its latency significantly. Specifically, the latency is reduced by $79.3\%$ on A-847 and $77.0\%$ on A-150.
TABLE VII: Effect of Top $k$ categories on CRM.
Top $k$ categories A-847 PC-459 A-150 PC-59 PAS-20
1 13.8 20.5 35.8 56.9 95.4
3 13.9 20.5 35.8 56.9 95.4
5 13.8 20.5 35.8 56.8 95.4
TABLE VIII: The impact of Spatial-level Fusion Module (SFM) and Class-level Fusion Module (CFM).
ESCF A-847 PC-459 A-150 PC-59 PAS-20
SFM 13.1 19.2 34.4 56.3 94.4
CFM 13.0 19.6 34.1 55.4 95.5
Both 13.9 20.5 35.8 56.9 95.4
TABLE IX: Impact of star operation.
ESCF A-847 PC-459 A-150 PC-59 PAS-20
w/o star 14.1 20.1 35.6 56.3 94.7
with star 13.9 20.5 35.8 56.9 95.4
TABLE X: Effect of the Spatial-level Sequence Reduction (SSR) and Class-level Sequence Reduction (CSR).
SSR CSR A-847 PC-459 A-150 PC-59 GFLOPs Memory
13.5 20.0 35.4 57.0 188.1 9.5
14.2 20.0 35.9 56.9 148.9 7.8
14.2 19.9 35.6 57.0 134.6 5.9
13.9 20.5 35.8 56.9 95.3 4.2
IV-C Ablation Study
IV-C1 Component analysis
The ERR-Seg framework comprises four main branches: Hierarchical Semantic Module (HSM), Channel Reduction Module (CRM), Efficient Semantic Context Fusion (ESCF), and Gradual Upsampling Decoder (GUD). To assess the importance of these modules, each module is individually removed, and the results are shown in Table V. The findings demonstrate the contribution of ESCF and GUD in enhancing performance, with improvements of +$9.7$ and +$1.0$ mIoU on A-150, respectively. A more detailed efficiency analysis of the ESCF will be provided later. Without using CRM, the redundancy reduction process is eliminated during training and evaluation, resulting in a decrease of $0.4$ mIoU on PC-459 and $0.5$ mIoU on A-150. Additionally, HSM yields a +$1.2$ mIoU enhancement on PC-459 and a +$0.5$ mIoU improvement on A-150, indicating the significance of hierarchical cost maps.
IV-C2 Ablation study for HSM
Without HSM, the gradient for updating stage 4 of CLIP’s visual encoder is derived solely from $\mathcal{F}_{4}^{s}$. However, when HSM is active, the gradient can be sourced from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$, as HSM functions as a parallel branch. Table VI investigates the impact of different gradient sources on updating stage 4. The results clearly show that the combination of gradients from $\mathcal{F}_{3}^{s}$ and $\mathcal{F}_{4}^{s}$ yields superior outcomes compared to utilizing either source individually. Furthermore, relying solely on $\mathcal{F}_{3}^{s}$ leads to notably poorer model performance, since $\mathcal{F}_{4}^{s}$ represents the original CLIP output and offers more precise semantic details.
Figure 8: Visualization of segmentation results on ADE20K-150 and ADE20K-847. The results of previous state-of-the-art methods SAN [15] and CAT-Seg [16] are also included for comparison.
Figure 9: Visualization of segmentation results in various domains. Our proposed ERR-Seg is capable of segmenting capybaras (a rare category in public datasets) from various domains, including (a) synthesized images, (b) cartoon images, (c) natural images, and (d) capybara dolls. Moreover, ERR-Seg achieves more precise masks than SAN [14] and CAT-Seg [16].
IV-C3 Impact of different design of CRM
This section investigated the impact of hyperparameters within CRM. Fig. 6 illustrates how the score coefficient $\lambda$ influences class scoring, with optimal performance achieved at $\lambda=0.1$. A larger $\lambda$ may blur the distinction between classes with similar rankings, while a smaller $\lambda$ might overlook certain classes. Furthermore, Fig. 7 explores the effect of the selected class number $P$ during training on the COCO-Stuff dataset, highlighting that setting $P=32$ results in the highest performance. It also reveals that setting $P$ either too large or too small can hinder the modeling of contextual information. Table VII presents the effect of the top $k$ categories of a pixel, indicating that the CRM is minimally sensitive to $k$, with slightly superior results observed at $k=3$. A smaller $k$, such as 1, risks overlooking existing categories due to potential inaccuracies in cost maps’ prior knowledge. Conversely, a larger $k$ may introduce redundant categories, hindering CRM’s ability to identify the most relevant classes.
IV-C4 Ablation study for ESCF
ESCF aims to incorporate more contextual information at both the class and spatial levels within the cost maps. We observed a significant decrease in performance when the Spatial-level Fusion Module (SFM) and the Class-level Fusion Module (CFM) were removed, as shown in Table VIII. Additionally, Table IX presents the results obtained without the star operation but MLP instead. It shows that the performance of ERR-Seg is improved on four datasets through the star operation. Table X presents the effects of the sequence reduction strategies, SSR and CSR, including their impact on GFLOPs and memory usage (GiB). The findings indicate that the use of CSR or SSR alone does not negatively affect performance. Overall, ESCF achieves a reduction of $49.6\%$ GFLOPs and $55.6\%$ memory without compromising performance.
TABLE XI: Ablation study for GUD
Guidance Feature A-847 PC-459 A-150 PC-59 PAS-20
$\mathcal{F}_{1}^{v},\mathcal{F}_{2}^{v}$ 14.0 20.3 35.3 57.1 95.3
$\mathcal{F}_{1}^{v},\mathcal{F}_{2}^{v},\mathcal{F}_{3}^{v}$ 13.9 20.5 35.8 56.9 95.4
Guidance Gradient A-847 PC-459 A-150 PC-59 PAS-20
w/o stop 13.9 19.7 35.1 56.9 94.9
with stop 13.9 20.5 35.8 56.9 95.4
Kernel Size A-847 PC-459 A-150 PC-59 PAS-20
3 13.9 20.5 35.8 56.9 95.4
5 13.8 20.3 35.4 56.6 95.2
7 13.7 20.1 35.7 56.7 95.1
IV-C5 Ablation study for GUD
GUD utilizes the middle-layer features ($\mathcal{F}_{i}^{v},1\leq i\leq 3$) from the CLIP’s visual encoder to reconstruct high-resolution segmentation results. We first investigate the necessity of integrating $\mathcal{F}_{3}^{v}$, as it shares the same size as the original cost maps. The experimental results are shown in the top section of Table XI, demonstrating the effectiveness of incorporating with $\mathcal{F}_{3}^{v}$, resulting in a +$0.5$ mIoU improvement on A-150. Moreover, we investigate the impact of updating the visual encoder with the gradient of $\mathcal{F}_{i}^{v}(1\leq i\leq 3)$, as this could potentially disrupt the alignment between the visual and text encoders. The results are presented in the middle section of Table XI. It shows that preventing the gradient of $\mathcal{F}_{i}^{v}(1\leq i\leq 3)$ from updating the visual encoder leads to better performance. The lower section of Table XI presents the effect of convolution kernel size, revealing that a smaller kernel size improves performance due to the localized nature of textual information.
TABLE XII: Impact of CLIP Architecture
CLIP Arch. A-847 PC-459 A-150 PC-59 PAS-20
ViT-B/16 7.5 13.0 22.8 49.7 91.0
ConvNeXt-B 9.7 12.5 26.9 49.4 89.9
TABLE XIII: The impact of sampled class number of CRM during inference.
A-847 PC-459 A-150 PC-59 PAS-20
[…]
32 13.6 16 18.8 16 34.9 8 54.9 4 90.2
48 13.9 32 19.9 24 35.5 16 56.9 8 95.3
64 13.9 48 20.5 32 35.8 24 56.9 12 95.4
80 13.9 64 20.5 40 35.8 32 57.0 16 95.4
IV-C6 Impact of CLIP architecture
Following previous methods [15, 20], our ERR-Seg also adopts the hierarchical CLIP (ConvNeXt) instead of the plain CLIP (ViT). The impact of the CLIP architecture is illustrated in Table XII. To ensure a fair comparison, training and evaluation are conducted using an input size of $384^{2}$ with frozen CLIP. The results show that employing hierarchical CLIP leads to a +$2.2$ +and +$4.1$ mIoU improvement on A-847 and A-150, respectively. In comparison to plain CLIP, hierarchical CLIP offers the following advantages: (1) it allows the use of middle-layer features to enhance the reconstruction of high-resolution semantic maps, and (2) it permits direct inference on high-resolution images, eliminating the need for the sliding window inference strategy [16] employed in plain CLIP.
IV-C7 Impact of sampled class number
Table XIII illustrates the influence of the sampled class number $P$ in the CRM during the inference phase. It is evident that as $P$ increases gradually, the performance tends to reach a saturation point. This finding aligns with the expectation that a sufficient increase in $P$ can adequately cover all existing categories. Specifically, we assign the following values for $P$: 48 for A-847, 48 for PC-459, 32 for A-150, 24 for PC-59, and 16 for PAS-20.
IV-C8 Impact of input size
Table XIV presents the results with varying input sizes during the inference phase. The findings highlight a decrease in performance when input sizes are too large or too small. Specifically, for models based on ConvNeXt-B and ConvNeXt-L, the recommended input sizes are $640^{2}$ and $768^{2}$, respectively. Our analysis reveals that using ConvNeXt-L as the backbone of CLIP, there is a better alignment between the image and text spaces, enabling the extraction of more refined pixel-text cost maps and thus supporting larger input sizes.
TABLE XIV: The impact of crop size during inference.
Backbone Size A-847 PC-459 A-150 PC-59 PAS-20
ConvNeXt-B $512^{2}$ 13.7 19.8 35.1 56.7 95.3
$640^{2}$ 13.9 20.5 35.8 56.9 95.4
$768^{2}$ 13.5 20.1 35.7 57.1 95.1
$1024^{2}$ 13.1 18.7 34.3 55.5 93.3
ConvNeXt-L $512^{2}$ 16.0 22.9 36.9 60.0 96.7
$640^{2}$ 16.6 23.5 37.8 60.5 96.8
$768^{2}$ 16.9 23.9 38.3 60.7 96.6
$1024^{2}$ 16.6 23.4 38.2 59.8 96.2
Figure 10: Visualization of ERR-Seg’s attribute understanding. ERR-Seg can correctly distinguish between a yellow dog and a white dog and between a lying capybara and a standing capybara.
IV-D Qualitative Results
Fig. 8 presents a qualitative comparison among SAN [14], CAT-Seg [16], and our proposed ERR-Seg on the validation sets of ADE20K-150 and ADE20K-847. The results highlight the superiority of our method over previous approaches in mask generation and open-vocabulary recognition. In addition, ERR-Seg can identify targets that were previously overlooked in ADE20K. While the current OVSS model shows remarkable open-vocabulary recognition capability due to the advancements of VLMs, it struggles with segmenting intricate masks. There is still potential for improvement in effectively recognizing small and nested targets.
As presented in Fig. 9, ERR-Seg shows its impressive ability to segment and identify capybaras in various domains, highlighting its universal segmentation capability. The selection of capybaras is due to their rarity in public datasets. Furthermore, compared with SAN and CAT-Seg, ERR-Seg achieves more precise masks with reduced noise.
Fig. 10 presents two simple examples that illustrate the ability of ERR-Seg to understand attributes such as color and posture. Our model accurately segments dogs of various colors and capybaras in different postures, demonstrating its exceptional open-world understanding capabilities.
V Conclusion
In conclusion, we introduce ERR-Seg, a novel framework for open-vocabulary semantic segmentation (OVSS). ERR-Seg investigates the impact of redundant classes and reveals that they lead to long latency and hinder the attention mechanism from capturing spatial-level and class-level dependencies. To address this issue, ERR-Seg proposes the training-free Channel Reduction Module (CRM) to eliminate redundant classes based on prior knowledge from VLMs. Additionally, recognizing the importance of hierarchical semantics for semantic segmentation, ERR-Seg introduces the Hierarchical Semantic Module (HSM) to generate hierarchical cost maps from the open-vocabulary visual encoder. Experimental results demonstrate that ERR-Seg establishes a new state-of-the-art in terms of efficiency and accuracy. Future work will focus on exploring high-quality real-time OVSS architecture design.